OpenStudy (anonymous):

Solve the equation for y. 5x – 7y = 10 @mikaylabean @ganeshie8

OpenStudy (anonymous):

-y=10-5x/7

OpenStudy (anonymous):

that's not one of my choices

OpenStudy (anonymous):

Hm. y = 5/7x -10/7

OpenStudy (anonymous):

Is that a choice..?

OpenStudy (anonymous):

yes can you help me with some more and are u shure that's right?

OpenStudy (anonymous):

Positive.

OpenStudy (anonymous):

And yes

OpenStudy (anonymous):

Would y=5/7x-10/7 not be correct?

OpenStudy (anonymous):

i see where @mikaylabean answered it!

OpenStudy (anonymous):

Which ordered pairs in the form (x, y) are solutions to the equation 3x – 4y = 21? Choose all answers that are correct. A. (−3, 3) B. (−1, −6) C. (7, 0) D. (11, 3)

OpenStudy (anonymous):

B!

OpenStudy (anonymous):

Which of the ordered pairs in the form (x, y) is a solution of this equation? A. The first is a solution, but the second is not. B. Neither is a solution. C. Both are solutions. D. The first is not a solution, but the second is